OpenStudy (anonymous):

Find the formula for the given two series: {1,2,4,5,7,8,10,11,13,14,16,17,19,20,22...} {1,2,4,5,8,9,13,14,19,20,26,27,34,35,43...} 1st series = +1, +2, +1, +2, +1, +2... 2nd series = +1, +2, +1, +3, +1, +4...

OpenStudy (rizags):

formula as in.... like arithmetic sequence formula?

OpenStudy (anonymous):

No, the difference between the terms aren't constant

OpenStudy (rizags):

right, but we can still write a formula

OpenStudy (anonymous):

Yes, but heavens if I can find it

OpenStudy (rizags):

wait i almost got it

OpenStudy (rizags):

nope nevermind sorry

OpenStudy (anonymous):

You can define the associated sequences recursively. Not so sure about a closed formula for the \(n\)-th term though... For the first sequence, you could say \[\begin{cases}a_n=\begin{cases}a_{n-1}+2&\text{for odd }n\\a_{n-1}+1&\text{for even }n\end{cases}\\a_1=1\end{cases}\] For the second, \[\begin{cases}a_n=\begin{cases}a_{n-1}+\left\lceil\dfrac{\color{red}{n+1}}{2}\right\rceil&\text{for odd }n\\a_{n-1}+1&\text{for even }n\end{cases}\\a_1=1\end{cases}\] (\(n+1\) can also be replaced with \(n+2\)).